OpenStudy (anonymous):

4√18 = 12√3? srry i meant 18 not 8

OpenStudy (anonymous):

Nope

jimthompson5910 (jim_thompson5910):

\[\Large 4\sqrt{18} \ \overset{?}{=}12\sqrt{3}\] \[\Large 4\sqrt{9*2} \ \overset{?}{=}12\sqrt{3}\] \[\Large 4\sqrt{9}*\sqrt{2} \ \overset{?}{=}12\sqrt{3}\] \[\Large 4*3*\sqrt{2} \ \overset{?}{=}12\sqrt{3}\] \[\Large 12\sqrt{2} \ \neq12\sqrt{3}\] So \[\Large 4\sqrt{18} \neq 12\sqrt{3}\]